A small town library buys a total of 32 books that cost $458. Some of the books cost $13 each, and the remainder cost $15 each. How many books of each price did they buy?
Answer by SHUgrad05(58)

Let x= the # books that cost $13; and y= the # books that cost $15.
Set up a system of equations:
x+y=32
13x+15y=458
Solve by subsitution:
x+y=32
y=32-x (plug into second equation)
13x+15(32-x)=458
13x+480-15x=458 -> move x to one side
-2x=-22
x=11 -> (plug into first equation)
11+y=32
y=21
There are 11 books at $13; 21 books at $15
Check:
11*$13=$143
21*$15=$315
=$458